OpenStudy (anonymous):

When submitting an answer in the form of an exponent, use a caret (^) to indicate the power. For example, 53 would be written as 5^3. Factor each expression. 1. 216x3 – 1 (1 point) 2. 8x3 + 125 (1 point) 3. x4 – 5x2 + 4 (1 point)

OpenStudy (anonymous):

it turns out that \(216=6^3\) so the first one is the difference of two cubes \[(6x)^3-1^3\] use \[a^3-b^3=(a-b)(a^2+ab+b^2)\] with \(a=6x\) and \(b=1\)

OpenStudy (perl):

Prove that if there is an injection f : A -> A which is not surjective, then A is not a finite set.

OpenStudy (perl):

satellite, maybe you can help me

OpenStudy (perl):

i have heard you are a math genius

OpenStudy (anonymous):

similarly, the second one is the sum of two cubes, \[(2x)^3+5^3\] this time use \[a^3+b^3=(a+b)(a^2-ab+b^2)\] with \[a=2x,b-5\]

OpenStudy (anonymous):

@perl post an i will see what we can do

OpenStudy (perl):

ok :)

OpenStudy (perl):

satellite, you do statistics?

OpenStudy (anonymous):

thank you sattelite